OpenStudy (anonymous):

You must show your work on each of the following questions. 1. Rewrite y = x^2 + 14x + 29 in general form. 2. Rewrite y = 3x^2 - 24x + 10 in general form. 3. Solve for x: (x - 9)^2 = 1 4. Solve for x: x^2 + 24x + 90 = 0 5. Solve for x: 2x^2 - 4x - 14 = 0 6. Create your own quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square.

OpenStudy (anonymous):

1....already in general form. 2. ...already in general form

OpenStudy (anonymous):

3)\[(x-9)^2=1\] \[(\sqrt{(x-9)^2}=\sqrt{1}\] x-9=1 x=1+9,x=10

OpenStudy (anonymous):

4) the general solution using the quadratic equation is: \[x=\frac{ -24\pm \sqrt{(2-24)^2-(4*1*90)} }{ 2*1 }=\frac{ -24\pm \sqrt{216} }{ 2 }\] \[x=\frac{ -24\pm6\sqrt{6} }{ 2 }\] \[x=-12\pm3\sqrt{6}\]

OpenStudy (anonymous):

5) the general solution using the quadratic equation is: \[x=\frac{ 4\pm \sqrt{(-4)^2-(4*2(-14))} }{2*2 }=\frac{ 4\pm \sqrt{128} }{ 4}\] \[x=1\pm2\sqrt{2}\]

OpenStudy (anonymous):

6) x^2 - 6x = 0 by graphing:

OpenStudy (anonymous):

x=6,x=0

OpenStudy (anonymous):

by factoring: x(x-6)=0 x=0 or x-6=0 x=6

OpenStudy (anonymous):

by the quadratic formula: \[x=\frac{ 6\pm \sqrt{(-6)^2-(4*1*0)} }{2*1}=\frac{ 6\pm6 }{ 2 }\] x=0,x=6

OpenStudy (anonymous):

by completing the square: x^2 - 6x +(-6/2)^2= (-6/2)^2 x^2-6x+(-6/2)^2=9 (x+(-6/2))^2=9 (x-3)^2=9 x-3=3,x=6 or x-3=-3,x=0

OpenStudy (anonymous):

@touseii45 get it.

OpenStudy (anonymous):

I get the most of it (: thank you so very much! @ASAAD123

OpenStudy (anonymous):

welcome ^-^.